# White Men Lead, Black Women Help? Benchmarking Language Agency Social Biases in LLMs

### v3

@@ 4.2 Findings 2: LLMs Suffer From Gender, Racial, and Especially Intersectional Biases in Language Agency
Table  3 demonstrates full results for gender, racial, and intersectional biases in language agency for biographies, professor reviews, and reference letters generated by the investigated 3 LLMs. We also visualize the average agentic-communal ratio gap in texts describing different gender and racial intersectional groups as overlapping horizontal bar graphs in Figure 4.
In the gender bias dimension, LLMs tend to depict males with more agentic language than females. As discussed in Section  4.1, all 3 LLMs possess notable levels of gender differences in agentic-communal ratio gaps. Table  3 further shows high variances of agency levels across gender groups. Both observations reveal notable language agency gender biases in LLM-generated texts.
In the racial bias dimension, LLM-generated texts for colored individuals are often remarkably less agentic than those for White individuals. Across all generation tasks, LLM-written texts about colored individuals have notably lower agency level than those for White individuals. For instance, as shown in Figure  3, Black professors receive reviews with the lowest agency levels in Chatgpt- and Llama3-generated reviews; huge discrepancies can be observed between agentic-communal ratio gaps in reviews for Black faculties and for professors of other races. Interestingly, studies on real-world professor ratings also found that Black professors received more negative reviews from students (Reid, 2010). Similarly, LLM-generated reference letters for White individuals are highest in agency, whereas Black individuals receive letters with the lowest language agency level, aligning with previous social science findings on racial biases (Powers et al., 2020; Chapman et al., 2022).
In intersectional bias dimension, texts depicting individuals at the intersection of gender and racial minority groups—such as Black females—possess remarkably lower language agency levels. Both quantitative results in Appendix E Tables 20, 22,24 and visualized illustrations in Figure  3 show severe intersectional biases across all LLMs on all generation tasks—those who are at the intersection of gender and racial minority groups are the most vulnerable to biases in language agency. For instance, ChatGPT- and Llama3-generated reviews for Black female professors show the lowest level of agency across all intersectional groups. Interestingly, we observe that on all text generation tasks, language agency is notably higher in texts about males within each racial group (e.g. Black males are described with more agentic language than Black females). These observations further align with prior social science findings on intersectional biases targeting gender and racial minority groups in texts (10.1145/3485447.3512134; Adams et al., 2019; Lemieux et al., 2023; Otterbacher, 2015; Chávez and Mitchell, 2020).
@@ 4.3 Findings 3: All LLMs Show Agency Biases, and Llama3 is the Most Biased One
Quantitative results in Table 3 show that prompt-based methods fail to effectively resolve language agency bias. More shockingly, we reveal that fairness instructions in prompts could even result in higher bias levels in LLM-generated texts. We found 2 main problems with prompt-based mitigation methods in LLMs: (1) bias amplification, for which even more severe biases are observed, and (2) bias overshooting, where biases are shifted towards anti-stereotype directions (e.g. biased towards males). In Table  4, we provide quantitative examples for these problems for the 3 LLMs.
@@ Appendix A Prompt Construction for Text Generation Tasks in LABE
We adapt the Context-Less Generation (CLG) approach proposed by  Wan et al. [2023] to prompt LLMs on biography, professor review, and reference letter generation tasks using descriptor-based prompt templates. All prompts contain gender and race descriptors to include demographic information. Additional information such as name, occupation, age, and department are also injected in prompts for different tasks. Table 5 demonstrates prompt templates used in LABE to obtain LLM outputs on the 3 text generation tasks about people with different demographic identities, for comprehensive bias analysis. Table 7 presents lists of descriptors including gender, race, occupations, departments, and ages. To identify natural names for males and females in different racial groups, we prompted ChatGPT to generate $5$ popular male names and $5$ popular female names in each of the $4$ race categories: “White”, “Black”, “Hispanic”, and “Asian”. Table 6 shows the full list of investigated racial groups and corresponding popular names for males and females within each group.
@@ B.1 Preprocessing
For the base dataset, we utilize the “evaluation” split of WikiBio Lebret et al. [2016], a personal biography dataset with information extracted from Wikipedia. We preprocess the dataset by splitting each personal biography into sentences. To ensure that each sentence is informative and depicts the owner of the biography, we remove the first two sentences and the last sentence, which usually provide the birth date and the current status of the owners without describing their characteristics.
@@ B.2 ChatGPT Generation
For each of the pre-processed sentences in personal biographies, we prompt the gpt-3.5-turbo-1106 version of ChatGPT with one-shot example Wang et al. [2020] to paraphrase it into an agentic version and a communal version. Specific prompt used in the dataset generation process is provided in Table 8. This guarantees the balance of the constructed dataset and prevents the propagation of pre-existing biases in the classifier training process.
@@ B.4 Post-processing
Appendix C Building Language Agency Classifiers with LAC
We provide details of training and inferencing language agency classification tools below. Licensing information for each model involved is provided in Table 11.
C.1 Model Choices
We experiment with BERT and RoBERTa to build discriminative classifiers for language agency. For generative classifier, we experiment with the Reinforcement Learning with Human Feedback (RLHF)-tuned Llama2 for dialogue use cases Touvron et al. [2023]. Below, we provide details on training and inferencing the models. For BERT and RoBERTa, we conduct full fine-tuning. For Llama2, we test with zero-shot prompting, one-shot prompting, and LoRA fine-tuning.
@@ C.1 Model Choices
Generative Model For the Llama2 generative model, we experiment with $4$ different settings: zero-shot prompting without definition, zero-shot prompting with definition, one-shot prompting with definition and an example, and parameter-efficient fine-tuning with LoRA Hu et al. [2021]. For reproducibility, we provide the full prompts used to probe Llama2 in zero-shot and few-shot settings in Table 10. For LoRA fine-tuning, we use a learning rate of $5e^{-5}$ to train for $5$ epochs. During inference, we follow the default generation configuration to set top-p to $1.0$, tok-k to $50$, and temperature to $1.0$.
@@
[Wan et al., 2023] 109M MIT License + Fine-Tune $66.49$ $66.49$ $64.22$ $64.82$
@@
C.2 Model Performance
We report the performances of baseline methods to classify language agency, as well as our trained classifiers on the LAC dataset. For baseline methods, we experimented on string matching, sentiment classification, and the agency classifier proposed in Wan et al. [2023]’s work. For string matching, we utilized Stahl et al. [2022]’s released lists of agentic and communal words with no licensing information. For sentiment classification, we utilized the sentiment classification pipeline in the transformers library with the off-the-shelf “distilbert/distilbert-base-uncased-finetuned-sst-2-english” [^3] https://huggingface.co/distilbert/distilbert-base-uncased-finetuned-sst-2-english model.
@@ C.2 Model Performance
Appendix D Human-Written Datasets Details
@@
Bias in Bios   The Bias in Bios De-Arteaga et al. [2019] dataset is released under MIT license. For preprocessing this dataset, we randomly sample $120$ biographies for each gender for each of the professions. Table 12 shows the full list of professions in the pre-processed dataset.
@@
Reference Letter [Wan et al., 2023] We directly utilize the reference letter dataset generated under the CBG setting in  Wan et al. [2023]’s work, which is released under the MIT License.
Appendix E Additional Experiment Results
We hereby provide additional experiment results on (1) stratified analysis on the Bias in Bios and RateMyProfessor Dataset, and (2) full evaluation results across the 3 LLMs, 3 text generation tasks, and all investigated gender, racial, and intersectional demographic groups.
E.1 Stratified Analysis on Human-Written Datasets
We stratify analysis on the human-written biography dataset based on professions and provide full results in Table 15. We then visualize the top $8$ most biased occupations as overlap horizontal bar graphs in Figure 4. Drastic language agency gender biases are found for pastor, architect, and software engineer. Interestingly, real-world reports have also demonstrated male dominance and gender bias in these occupations [Kathleen Schubring, , A. Nicholson et al., , Kaminski, ]. Similarly, we stratify our analysis on the human-written professor review dataset based on academic departments in Table 14, and visualize the top $8$ most biased departments in Figure 4. Greatest biases are observed in reviews for professors in departments such as Accounting, Sociology, and Chemistry; all $3$ departments have been proven to be male-dominated [200, 2009, Girgus, , Seijo, ]. Language agency gender biases found on human-written texts in our study align with findings of social science studies, showing that our proposed evaluation tools effectively capture implicit language style biases.
@@ E.1 Stratified Analysis on Human-Written Datasets
E.2 Full Evaluation Results
@@
Table  16 shows results for gender biases before mitigation, whereas Table  17 presents results after mitigation. Table  18 presents results for racial biases before mitigation, and Table  19 shows results after mitigation. For intersectional biases, results for ChatGPT before mitigation are in Table  20; results after mitigation are in Table  21. Intersectional results for Mistral before mitigation are in Table  22; results after mitigation are in Table  23. Intersectional outcomes for Llama3 before mitigation are in Table  24; results after mitigation are in Table  25.
@@
Reference Letter [Wan et al., 2023] Male 57.47 42.53 14.94 4.64
@@
Appendix F Computational Resources
Appendix G Limitations
We identify some limitations of our study. First, due to the limited information within the datasets available for our study, we were only able to consider the binary gender and $4$ racial groups for bias analyses. However, we note that it is important and significant for further works to extend the investigation of the fairness problem in our study to other gender and racial minority groups. Second, due to the scarcity of data, our study were only able to investigate language agency-related gender biases in 2 human-written datasets of personal biographies and professor reviews. We encourage future studies to extend the exploration of racial and intersectional language agency biases in broader domains of human-written texts. Third,, due to cost and resource constraints, we were not able to further extend our experiments to larger scales. Future works should be devoted to comprehensively evaluating biases from various data sources. Lastly, experiments in this study incorporate language models that were pre-trained on a wide range of text from the internet and have been shown to learn or amplify biases from the data used. Since we utilize a language model to synthesize a language agency classification dataset, we adopt a number of methods to prevent potential harm and bias propagation: (1) we prompt the model to paraphrase each input into an agentic version and a communal version, ensuring the balance in the preliminary generated dataset, and (2) we invite expert annotators to re-annotate the generated data, to verify and ensure the quality of the final dataset used to train language agency classifiers. Although these methods might not guarantee complete fairness, it is the best we can do to prevent bias propagation. We encourage future extensions of our works to also consider this factor in their research, so as to draw reliable and trustworthy research conclusions.

### v4

@@ 4.2 Findings 2: LLMs Suffer From Gender, Racial, and Especially Intersectional Biases in Language Agency
Table  1 demonstrates full results for gender, racial, and intersectional biases in language agency for biographies, professor reviews, and reference letters generated by the investigated 3 LLMs. We also visualize the average agentic-communal ratio gap in texts describing different gender and racial intersectional groups as overlapping horizontal bar graphs in Figure 4.
In the gender bias dimension, LLMs tend to depict males with more agentic language than females. As discussed in Section  4.1, all 3 LLMs possess notable levels of gender differences in agentic-communal ratio gaps. Table  1 further shows high variances of agency levels across gender groups. Both observations reveal notable language agency gender biases in LLM-generated texts.
@@
In the racial bias dimension, LLM-generated texts for colored individuals are often remarkably less agentic than those for White individuals. Across all generation tasks, LLM-written texts about colored individuals have notably lower agency level than those for White individuals. For instance, as shown in Figure  3, Black professors receive reviews with the lowest agency levels in Chatgpt- and Llama3-generated reviews; huge discrepancies can be observed between agentic-communal ratio gaps in reviews for Black faculties and for professors of other races. Interestingly, studies on real-world professor ratings also found that Black professors received more negative reviews from students (Reid, 2010). Similarly, LLM-generated reference letters for White individuals are highest in agency, whereas those for Black individuals have the lowest language agency, aligning with previous social science findings on racial biases (Powers et al., 2020; Chapman et al., 2022).
In intersectional bias dimension, texts depicting individuals at the intersection of gender and racial minority groups—such as Black females—possess remarkably lower language agency levels. Both quantitative results in Appendix D Tables 21, 23,25 and visualized illustrations in Figure  3 show severe intersectional biases across all LLMs on all generation tasks—those who are at the intersection of gender and racial minority groups are the most vulnerable to biases in language agency. For instance, ChatGPT- and Llama3-generated reviews for Black female professors show the lowest level of agency across all intersectional groups. Interestingly, we observe that on all text generation tasks, language agency is notably higher in texts about males within each racial group (e.g. Black males are described with more agentic language than Black females). These observations further align with prior social science findings on intersectional biases targeting gender and racial minority groups in texts (Field et al., 2022; Adams et al., 2019; Lemieux et al., 2023; Otterbacher, 2015; Chávez and Mitchell, 2020).
@@ 4.3 Findings 3: All LLMs Show Agency Biases, and Llama3 is the Most Biased One
Quantitative results in Table 1 show that prompt-based methods fail to effectively resolve language agency bias. More shockingly, we reveal that fairness instructions in prompts could even result in higher bias levels in LLM-generated texts. We found 2 main problems with prompt-based mitigation methods in LLMs: (1) bias amplification, for which even more severe biases are observed, and (2) bias overshooting, where biases are shifted towards anti-stereotype directions (e.g. biased towards males). In Table  2, we provide quantitative examples for these problems for the 3 LLMs.
@@ 5 Conclusion
Limitations
We identify some limitations of our study. First, due to the limited information within the datasets available for our study, we were only able to consider the binary gender and $4$ racial groups for bias analyses. However, we note that it is important and significant for further works to extend the investigation of the fairness problem in our study to other gender and racial minority groups. Second, due to the scarcity of data, our study were only able to investigate language agency-related gender biases in 2 human-written datasets of personal biographies and professor reviews. We encourage future studies to extend the exploration of racial and intersectional language agency biases in broader domains of human-written texts. Third, due to cost and resource constraints, we were not able to further extend our experiments to larger scales. Future works should be devoted to comprehensively evaluating biases from various data sources. Lastly, experiments in this study incorporate language models that were pre-trained on a wide range of text from the internet and have been shown to learn or amplify biases from the data used. Since we utilize a language model to synthesize a language agency classification dataset, we adopt a number of methods to prevent potential harm and bias propagation: (1) we prompt the model to paraphrase each input into an agentic version and a communal version, ensuring the balance in the preliminary generated dataset, and (2) we invite expert annotators to re-annotate the generated data, to verify and ensure the quality of the final dataset used to train language agency classifiers. Although these methods might not guarantee complete fairness, it is the best we can do to prevent bias propagation. We encourage future extensions of our works to also consider this factor in their research, so as to draw reliable and trustworthy research conclusions.
Ethics Statement
Experiments in this study incorporate language models that were pre-trained on a wide range of text from the internet and have been shown to learn or amplify biases from the data used. Since we utilize a language model to synthesize a language agency classification dataset, we adopt a number of methods to prevent potential harm and bias propagation: (1) we prompt the model to paraphrase each input into an agentic version and a communal version, ensuring the balance in the preliminary generated dataset, and (2) we invite expert annotators to re-annotate the generated data, to verify and ensure the quality of the final dataset used to train language agency classifiers. We encourage future extensions of our works to also consider this factor in their research, so as to draw reliable and trustworthy research conclusions.
@@ Appendix A Prompt Construction for Text Generation Tasks in LABE
We adapt the Context-Less Generation (CLG) approach proposed by  Wan et al. (2023) to prompt LLMs on biography, professor review, and reference letter generation tasks using descriptor-based prompt templates. All prompts contain gender and race descriptors to include demographic information. Additional information such as name, occupation, age, and department are also injected in prompts for different tasks. Table 3 demonstrates prompt templates used in LABE to obtain LLM outputs on the 3 text generation tasks about people with different demographic identities, for comprehensive bias analysis. Table 5 presents lists of descriptors including gender, race, occupations, departments, and ages. To identify natural names for males and females in different racial groups, we prompted ChatGPT to generate $5$ popular male names and $5$ popular female names in each of the $4$ race categories: “White”, “Black”, “Hispanic”, and “Asian”. Table 4 shows the full list of investigated racial groups and corresponding popular names for males and females within each group.
@@ B.1 Preprocessing
For the base dataset, we utilize the “evaluation” split of WikiBio Lebret et al. (2016), a personal biography dataset with information extracted from Wikipedia. We preprocess the dataset by splitting each personal biography into sentences. To ensure that each sentence is informative and depicts the owner of the biography, we remove the first two sentences and the last sentence, which usually provide the birth date and the current status of the owners without describing their characteristics.
@@ B.2 ChatGPT Generation
For each of the pre-processed sentences in personal biographies, we prompt the gpt-3.5-turbo-1106 version of ChatGPT with one-shot example Wang et al. (2020) to paraphrase it into an agentic version and a communal version. Specific prompt used in the dataset generation process is provided in Table 6. This guarantees the balance of the constructed dataset and prevents the propagation of pre-existing biases in the classifier training process.
@@ B.4 Post-processing
B.5 Dataset Statistics
Split # Entries # Agentic # Communal
Train $2,979$ $1,523$ $1,456$
Valid $372$ $192$ $180$
Test $373$ $199$ $174$
Table 8: LAC Dataset Statistics.
Text Type Template # Entries
Table 9: Templates for LLM Generations in LABE.
B.6 Model Choices
We experiment with BERT and RoBERTa to build discriminative classifiers for language agency. For generative classifier, we experiment with the Reinforcement Learning with Human Feedback (RLHF)-tuned Llama2 for dialogue use cases Touvron et al. (2023). Below, we provide details on training and inferencing the models. For BERT and RoBERTa, we conduct full fine-tuning. For Llama2, we test with zero-shot prompting, one-shot prompting, and LoRA fine-tuning.
@@ B.6 Model Choices
Generative Model For the Llama2 generative model, we experiment with $4$ different settings: zero-shot prompting without definition, zero-shot prompting with definition, one-shot prompting with definition and an example, and parameter-efficient fine-tuning with LoRA Hu et al. (2021). For reproducibility, we provide the full prompts used to probe Llama2 in zero-shot and few-shot settings in Table 10. For LoRA fine-tuning, we use a learning rate of $5e^{-5}$ to train for $5$ epochs. During inference, we follow the default generation configuration to set top-p to $1.0$, tok-k to $50$, and temperature to $1.0$.
@@
(Wan et al., 2023) 109M MIT License + Fine-Tune $66.49$ $66.49$ $64.22$ $64.82$
@@
B.7 Model Performance
We report the performances of baseline methods to classify language agency, as well as our trained classifiers on the LAC dataset. For baseline methods, we experimented on string matching, sentiment classification, and the agency classifier proposed in Wan et al. (2023)’s work. For string matching, we utilized Stahl et al. (2022)’s released lists of agentic and communal words with no licensing information. For sentiment classification, we utilized the sentiment classification pipeline in the transformers library with the off-the-shelf “distilbert/distilbert-base-uncased-finetuned-sst-2-english” [^3] https://huggingface.co/distilbert/distilbert-base-uncased-finetuned-sst-2-english model.
@@ B.7 Model Performance
Appendix C Human-Written Datasets Details
@@
Bias in Bios   The Bias in Bios De-Arteaga et al. (2019) dataset is released under MIT license. For preprocessing this dataset, we randomly sample $120$ biographies for each gender for each of the professions. Table 12 shows the full list of professions in the pre-processed dataset.
@@
Reference Letter (Wan et al., 2023) We directly utilize the reference letter dataset generated under the CBG setting in  Wan et al. (2023)’s work, which is released under the MIT License.
Appendix D Additional Experiment Results
We hereby provide additional experiment results on: (1) quantitative results on gender bias in LLM- vs. human-written texts, (2) stratified analysis on the Bias in Bios and RateMyProfessor Dataset, and (3) full evaluation results across the 3 LLMs, 3 text generation tasks, and all investigated gender, racial, and intersectional demographic groups.
D.1 Quantitative Results: Bias in LLMs vs. Human-Written Texts
Table 14 presents quantitative results on gender bias in human- and LLM-generated texts. Figure 2 provides a visualization of results in the table.
Dataset Model Gender Diff. (M-F)
Biography Human $10.12$
ChatGPT $8.49$
Mistral 10.87
Llama3 $8.51$
Professor Review Human $1.86$
ChatGPT $6.57$
Mistral $8.14$
Llama3 $11.51$
Reference Letter Wan et al. (2023) $4.64$
ChatGPT $9.33$
Mistral $10.84$
Llama3 $9.44$
Table 14: Language agency gender bias in human-written and LLM-generated texts, measured by gender difference in agency-communal ratio gaps. Highest bias for each type of text is in bold.
D.2 Stratified Analysis on Human-Written Datasets
We stratify analysis on the human-written biography dataset based on professions and provide full results in Table 16. We then visualize the top $8$ most biased occupations as overlap horizontal bar graphs in Figure 4. Drastic language agency gender biases are found for pastor, architect, and software engineer. Interestingly, real-world reports have also demonstrated male dominance and gender bias in these occupations (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Similarly, we stratify our analysis on the human-written professor review dataset based on academic departments in Table 15, and visualize the top $8$ most biased departments in Figure 4. Greatest biases are observed in reviews for professors in departments such as Accounting, Sociology, and Chemistry; all $3$ departments have been proven to be male-dominated (200, 2009; Girgus, ; Seijo, ). Language agency gender biases found on human-written texts in our study align with findings of social science studies, showing that our proposed evaluation tools effectively capture implicit language style biases.
@@ D.2 Stratified Analysis on Human-Written Datasets
D.3 Full Evaluation Results
@@
Table  17 shows results for gender biases before mitigation, whereas Table  18 presents results after mitigation. Table  19 presents results for racial biases before mitigation, and Table  20 shows results after mitigation. For intersectional biases, results for ChatGPT before mitigation are in Table  21; results after mitigation are in Table  22. Intersectional results for Mistral before mitigation are in Table  23; results after mitigation are in Table  24. Intersectional outcomes for Llama3 before mitigation are in Table  25; results after mitigation are in Table  26.
Appendix E Computational Resources
@@
Reference Letter (Wan et al., 2023) Male 57.47 42.53 14.94 4.64